OpenStudy (anonymous):

10% of what number is 70?

OpenStudy (campbell_st):

well if 10% = 70 multiply each by 10 and you get the answer 100% =

OpenStudy (anonymous):

@campbell_st 700?

OpenStudy (anonymous):

yes

OpenStudy (campbell_st):

thats correct... 70*10 = 700

OpenStudy (anonymous):

here is an easy way to find using fractions:\[\frac{ 10 }{ 100 }=\frac{ 70 }{ x }\]

OpenStudy (anonymous):

@campbell_st ty!

OpenStudy (anonymous):

o wow juliette

OpenStudy (anonymous):

@spencej22 what?

OpenStudy (anonymous):

i answered you first.

OpenStudy (anonymous):

lol whatever

OpenStudy (anonymous):

@spencej22 yes but you searched it up and I know that because it didnt say you where typing (: good day

OpenStudy (anonymous):

no i didnt search it that is so easy i did it in my head

OpenStudy (anonymous):

@spencej22 mm sure.

OpenStudy (anonymous):

i was 1st to reply